Sodium only has one naturally occuring isotope, Na23 , with a relative atomic mass of 22.9898 u. A synthetic, radioactive isotope of sodium, Na22 , is used in positron emission tomography. Na22 has a relative atomic mass of 21.9944 u. 


A 1.3476 g sample of sodium containing a mixture of Na23 and Na22 has an apparent "atomic mass" of 22.9501 u . Find the mass of Na22 contained in this sample.


1
Expert's answer
2021-09-12T23:18:17-0400

Let X= percentage of Na22 in decimal


Therefore ( 1-X)= percentage of Na23 in decimal.


21.9944X+22.9898(1-X)=22.9501.


Calculating for X:


X= 0.039883464


Therefore mass of Na22 in the sample =1.3476×0.039883464


=0.053746956g